# supplementary

### v1

@@ Abstract
In reinforcement learning (RL), agents often struggle to perform well on tasks that differ from those encountered during training. This limitation presents a challenge to the broader deployment of RL in diverse and dynamic task settings. In this work, we introduce memory augmentation, a memory-based RL approach to improve task generalization. Our approach leverages task-structured augmentations to simulate plausible out-of-distribution scenarios and incorporates memory mechanisms to enable context-aware policy adaptation. Trained on a predefined set of tasks, our policy demonstrates the ability to generalize to unseen tasks through memory augmentation without requiring additional interactions with the environment. Through extensive simulation experiments and real-world hardware evaluations on legged locomotion tasks, we demonstrate that our approach achieves zero-shot generalization to unseen tasks while maintaining robust in-distribution performance and high sample efficiency.
11footnotetext: Department of Mechanical and Process Engineering, ETH Zurich, Switzerland.22footnotetext: ETH AI Center, ETH Zurich, Switzerland.33footnotetext: Department of Computer Science, ETH Zurich, Switzerland.
{kaibao, chenhli, yardas, krausea, mahutter}@ethz.ch
I Introduction
@@
Reinforcement learning (RL) has made significant strides in training agents to perform complex tasks [1, 2, 3, 4, 5], achieving high performance in controlled settings. However, RL agents often struggle when faced with tasks outside their training distribution—a limitation known as the out-of-distribution (OOD) generalization challenge. This limitation restricts the applicability of RL policies to a broader range of tasks beyond those explicitly encountered during training [6, 7].
Figure 1: Improving task generalization through memory augmentation. Top row: RL agents often struggle with tasks not encountered during training, leading to poor performance. Middle row: Experience augmentation expands the range of training scenarios to include potential test-time tasks; however, in partially observable settings, agents may fail to identify the current task parameterizations, resulting in over-conservative behavior. Bottom row: Incorporating memory mechanisms enables agents to infer task context, facilitating context-aware decision-making and improved adaptability.
@@
A common way to address this challenge is to increase training diversity to better encompass potential testing conditions. Domain randomization [8, 9, 10], for instance, achieves this by varying environment parameters, thereby exposing agents to a wide range of scenarios during training. However, as the extent of randomization increases, this method often suffers from reduced sample efficiency [11]. Additionally, increased variations during training can result in over-conservative strategies, especially in environments where these variations are not fully observable. In such settings, the agent may struggle to discern the current task parameterizations and only executes conservative moves, thereby sacrificing optimality. This raises a crucial question: how can an RL agent achieve robust performance across diverse environment variations, or tasks, without being explicitly exposed to them during training?
@@
A promising method to improve the generalization of the policy is through experience augmentation [12], where training experiences are augmented to simulate a broader range of potential test-time conditions. However, pure augmentation, without properly accounting for task context, may similarly cause over-conservativeness resulting from exposing the policy to excessive unobservable variations. Building on the contextual modeling capability of Recurrent Neural Networks (RNNs), which have been shown to effectively address partial observability [13], we enhance experience augmentation in RL by incorporating memory mechanisms that respect the necessary task context, as shown in Fig. 1.
@@
II Related work
II-A Task generalization in RL
In RL, task generalization arises as a challenge when the agent is required to perform well on tasks that differ from those encountered during training. These differences may stem from variations in scenes, states, or reward structures [14]. Domain randomization, originally introduced for sim-to-real transfer in robotics [8, 9, 10], addresses this issue by randomizing environment parameters during training to encompass a broader range of potential test-time conditions. In robotic control, this often involves dynamics randomization, where parameters like masses, friction coefficients, joint damping, and actuator gains are varied during training to enhance policy robustness [10, 1, 2, 3, 15]. While this effectively broadens the agent’s exposure to diverse conditions, increased randomization often leads to reduced sample efficiency [11].
Alternatively, many works incorporate known physics principles or state structures into task design to approach task generalization. Examples include foot-placement dynamics [16], object invariance [17], granular media interactions [18], frequency domain parameterization [7], rigid body dynamics [19], and semi-structured Lagrangian dynamics models [20].
II-B Experience augmentation
While many data augmentation techniques in RL have focused on augmenting image-based observations to improve generalization across visual changes [21, 22, 23, 24], fewer studies have investigated augmentation at the induced state-action trajectory level—referred to as experience augmentation—which is the focus of our work.
One example of experience augmentation is PlayVirtual [25], which augments cycle-consistent virtual trajectories by leveraging forward and backward dynamics models in a latent space to improve feature representation learning. Meanwhile, in HER [12], which aims to address the sparse reward problem in RL, failed trajectories are transformed into meaningful learning signals through goal relabeling. This relabeling mechanism, as implemented in HER and its extensions [26, 27, 28, 29], can also be viewed as a form of experience augmentation. However, these approaches primarily focus on improving learning efficiency with limited training data rather than explicitly addressing task generalization.
In legged locomotion control, symmetry-based augmentation has been applied to state-action trajectories to encourage symmetry-invariant behavior [30, 31]. Inspired by these works, our approach leverages the inherent task structure to augment state-action trajectories, generating synthetic experience of novel scenarios to enhance task generalization.
II-C Meta-RL
Meta-RL enables agents to quickly adapt to new tasks by leveraging experience gained during meta-training. Through exposure to multiple related tasks, the agent learns a context-aware policy [32, 33, 34, 35] or a good initialization [36, 37, 38], which allows it to adapt to novel, unseen tasks with minimal additional training [39]. In the field of memory-based meta-RL, RNNs have been widely employed for their ability to capture task context from past interactions, enabling zero-shot adaptation [32, 33]. Building on this framework, Nagabandi et al. [40] developed a dynamics model prior to enable rapid online adaptation in dynamic environments, while Zargarbashi et al. [41] trained a morphology-agnostic locomotion policy for legged robots.
III Preliminaries
III-A Partially observable Markov Decision Process and Meta-RL
@@
We model the problem as a Partially Observable Markov Decision Process (POMDP). We define a POMDP as the tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{O},\mathcal{A},p,r,\gamma,\rho_{0}\rangle,$ where $\mathcal{S}$ is the set of states, $\mathcal{O}$ is the set of observations, $\mathcal{A}$ is the set of actions, $p:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ is the state transition function, $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is the reward function, $\gamma$ is the discount factor, and $\rho_{0}:\mathcal{S}\to[0,1]$ is the initial state distribution.
In POMDPs, observations provide only partial information about the true underlying state, requiring the agent to reason about the task context. In conventional RL, the agent aims to learn a policy $\pi$ that maximizes the expected cumulative reward:
@@
where $\tau=(s_{0},a_{0},s_{1},\dots)$ represents a trajectory sampled from the POMDP $\mathcal{M}$ under the policy $\pi$.
In this work, we frame the problem of learning across multiple tasks as a meta-RL problem, with the objective of optimizing over a distribution of tasks $p(\mathcal{T})$. In our setting, this corresponds to a distribution of POMDPs $\mathcal{M}(\mathcal{T})$. We focus on the scenario where each task $T\in\mathcal{T}$ exhibits unique dynamics while sharing a common reward function. The goal is to learn a policy $\pi$ that maximizes the expected cumulative reward across tasks:
@@
III-B Proximal Policy Optimization
In this work, we apply Proximal Policy Optimization (PPO) [42] as the RL algorithm. At the $k$-th learning step, the agent collects trajectories by executing its policy, parameterized by $\theta_{k}$. To optimize the policy, PPO employs importance sampling to address the discrepancy between the current policy $\pi_{\theta}(a_{t}\mid s_{t})$ and the policy from the previous update $\pi_{\theta_{k}}(a_{t}\mid s_{t})$ used to collect the data. The policy gradient is expressed as:
$\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}_{\tau\sim p_{\pi_{\theta_{k}}}}\left[\sum_{t=0}^{\infty}\eta_{t}(\theta)A^{\pi_{\theta_{k}}}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\right],$ (1)
where $\eta_{t}(\theta)=\frac{p_{\pi_{\theta}}(s_{t})}{p_{\pi_{\theta_{k}}}(s_{t})}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{k}}(a_{t}|s_{t})}$ is the importance sampling weight, and $A^{\pi_{\theta_{k}}}=A^{\pi_{\theta_{k}}}(s_{t},a_{t})$ is the advantage function under the policy $\pi_{\theta_{k}}$. To stabilize training, PPO optimizes a clipped surrogate loss $L_{\text{clip}}(\theta)$ [42], which constrains the magnitude of policy updates. Additionally, Generalized Advantage Estimation (GAE) [43] is commonly used to compute the advantage function, balancing bias and variance for efficient learning.
III-C Problem statement
@@
During training, only partial observations are available for $\mathcal{T}_{\text{ID}}$. In this setting, the agent must infer the task context implicitly based on past observations and actions.
IV Approach
IV-A Task-structured augmentation
To augment the training experience, our approach leverages a set of transformation functions, denoted by $\mathcal{G}$, where each $g\in\mathcal{G}$ represents a transformation that reflects the underlying structure that we assume to exist in the task space. These transformations are used to generate synthetic experience in $\mathcal{T_{\text{OOD}}^{\text{aug}}}$ from the original data collected in $\mathcal{T_{\text{ID}}}$.
Formally, let $\tau=\{(o_{t},a_{t},o_{t+1},r_{t})\}_{t=0}^{T}$ represent a trajectory collected by the current policy $\pi_{\theta_{k}}$ within the training set $\mathcal{T_{\text{ID}}}$. Each transformation $g=(g_{o},g_{a})\in\mathcal{G}$ consists of functions $g_{o}:\mathcal{O}\rightarrow\mathcal{O}$ and $g_{a}:\mathcal{A}\rightarrow\mathcal{A}$, which map an original observation-action pair $(o,a)$ to a new pair $(o^{g},a^{g})=(g_{o}(o),g_{a}(a))$. We denote $s_{t}^{g}$ as the latent state corresponding to the transformed observation $o_{t}^{g}$. For all $g\in\mathcal{G},\,s\in\mathcal{S},\,a\in\mathcal{A}$, it is assumed that the following properties hold:
@@ IV-A Task-structured augmentation
    Transition probability invariance: The probability of reaching the next state from a given state-action pair remains identical in their respective tasks. Formally, this is expressed as:
    $T(s_{t+1}^{g}|s_{t}^{g},a_{t}^{g})=T(s_{t+1}|s_{t},a_{t}).$
@@
    Initial state distribution invariance: The initial state distribution remains invariant across original and transformed trajectories. Formally, this is expressed as:
    $p_{0}(s^{g})=p_{0}(s).$
@@
    Reward invariance: Each original state-action pair $(s_{t},a_{t})$ and its transformed counterpart $(s_{t}^{g},a_{t}^{g})$ receive the same rewards in their respective tasks. Formally, this is expressed as:
    $r(s_{t}^{g},a_{t}^{g})=r(s_{t},a_{t}).$
These properties ensure alignment between the original and transformed tasks, maintaining the coherence of task dynamics. The resulting augmented trajectory $\tau_{g}=\{(o_{t}^{g},a_{t}^{g},o_{t+1}^{g},r_{t})\}_{t=0}^{T}$ represents the agent’s simulated experience within an augmented task $T\in\mathcal{T}_{\text{OOD}}^{\text{aug}}$, generated without real environment interactions. The training dataset is then constructed as the union of the original and augmented trajectories:
$\mathcal{D}_{\text{train}}=\mathcal{D}_{\text{ID}}\cup\mathcal{D}_{\text{OOD}}^{\text{aug}}=\{\tau_{i}\}_{i=1}^{N}\cup\bigcup_{g\in\mathcal{G}}\{\tau_{i}^{g}\}_{i=1}^{N}.$
An example of such transformations in $\mathcal{G}$ includes symmetry transformations, which have been used in RL for manipulation tasks [27] and legged locomotion control [30, 31]. Specifically, for quadrupedal robots, symmetry transformations can simulate mirrored conditions (e.g., joint failures) across left-right and front-back body parts. Similarly, for humanoid robots, they can replicate mirrored conditions across left-right body parts. These transformations leverage the morphological symmetry inherent to legged robots and uphold the essential invariance properties (a)(c) [44].
IV-B Training with augmented data
Following the method of Mittal et al. [30], we approach data augmentation in PPO by constructing augmented policies $\pi_{\theta_{k}}^{g}$ such that $\pi_{\theta_{k}}^{g}(a^{g}|s^{g})=\pi_{\theta_{k}}(a|s),\forall g\in\mathcal{G},s\in\mathcal{S},a\in\mathcal{A}.$ Based on properties (a)(c), it can be shown that for a state-action pair $(s,a)$ and its transformed counterpart $(s^{g},a^{g})$, the following invariances hold [30]:
$A^{\pi_{\theta_{k}}^{g}}(s^{g},a^{g})=A^{\pi_{\theta_{k}}}(s,a)\text{, and }p_{\pi_{\theta_{k}}^{g}}(s^{g})=p_{\pi_{\theta_{k}}}(s).$ (2)
Using Eq. 2, the policy update rule in Eq. 1 becomes
$\nabla_{\theta}J(\pi_{\theta})$ $=\mathbb{E}_{\tau\sim p_{\pi_{\theta_{k}}}}\left[\sum_{t=0}^{\infty}\eta_{t}(\theta)A^{\pi_{\theta_{k}}}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\right]$
$+\sum_{g\in G}\mathbb{E}_{\tau\sim p_{\pi_{\theta_{k}}}}\left[\sum_{t=0}^{\infty}\eta_{t}^{g}(\theta)A^{\pi_{\theta_{k}}}\nabla_{\theta}\log\pi_{\theta}(a_{t}^{g}|s_{t}^{g})\right],$ (3)
with $\eta_{t}^{g}(\theta)=\frac{p_{\pi_{\theta}}(s_{t}^{g})}{p_{\pi_{\theta_{k}}}(s_{t})}\frac{\pi_{\theta}(a_{t}^{g}|s_{t}^{g})}{\pi_{\theta_{k}}(a_{t}|s_{t})}$, which retains the action probability of the original data. For further details, we refer readers to [30].
@@ IV-B Training with augmented data
This approach ensures balanced learning across both the original and augmented tasks, promoting equitable representation for each task during the policy update.
IV-C Memory augmentation

### v2

# Toward Task Generalization via Memory Augmentation in Meta-Reinforcement Learning

@@ Abstract
Agents trained via reinforcement learning (RL) often struggle to perform well on tasks that differ from those encountered during training. This limitation presents a challenge to the broader deployment of RL in diverse and dynamic task settings. In this work, we introduce memory augmentation, a memory-based RL approach to improve task generalization. Our approach leverages task-structured augmentations to simulate plausible out-of-distribution scenarios and incorporates memory mechanisms to enable context-aware policy adaptation. Trained on a predefined set of tasks, our policy demonstrates the ability to generalize to unseen tasks through memory augmentation without requiring additional interactions with the environment. Through extensive simulation experiments and real-world hardware evaluations on legged locomotion tasks, we demonstrate that our approach achieves zero-shot generalization to unseen tasks while maintaining robust in-distribution performance and high sample efficiency.
Keywords: Meta-Reinforcement Learning, Memory Augmentation
1 Introduction
@@
Reinforcement learning (RL) has made significant strides in training agents to perform complex tasks [1, 2, 3, 4, 5], achieving high performance in controlled settings. However, RL agents often struggle when faced with tasks outside their training distribution—a limitation known as the out-of-distribution (OOD) generalization challenge. This limitation restricts the applicability of RL policies to a broader range of tasks beyond those encountered during training [6, 7].
@@
A common way to address this challenge is to increase training diversity to better encompass potential test-time conditions. Domain randomization [8, 9, 10], for instance, achieves this by varying environment parameters, thereby exposing agents to a wider range of training scenarios. However, as the extent of randomization increases, this method often suffers from reduced sample efficiency [11]. Additionally, increased variations during training can result in over-conservative strategies, especially in environments where these variations are not fully observable. In such settings, the agent may struggle to discern the current task parameterizations and only executes conservative moves, thereby sacrificing optimality. This raises a crucial question: how can an RL agent achieve robust performance across diverse environment variations, or tasks, without being explicitly exposed to them during training?
@@
A promising method to achieve this is through experience augmentation [12], where training experiences are augmented to simulate a broader range of potential test-time conditions. However, pure augmentation, without properly accounting for task context, may similarly cause over-conservativeness resulting from exposing the policy to excessive unobservable variations. Building on the contextual modeling capability of Recurrent Neural Networks (RNNs), which have been shown to effectively address partial observability [13], we enhance experience augmentation in RL by incorporating memory mechanisms that respect the necessary task context.
@@
Supplementary videos for this work are available at https://sites.google.com/view/corl2025-ma/home.
2 Related Work
Task generalization remains a fundamental challenge in RL, as policies trained on a fixed set of tasks often fail to perform well under novel task conditions. A common strategy to address this challenge is domain randomization, originally introduced for sim-to-real transfer in robotics [8, 9, 10]. This approach randomizes environment parameters such as masses, friction coefficients, joint damping, and actuator gains during training to expose the agent to a wider range of conditions  [10, 1, 2, 3, 14]. While it improves policy robustness, increased randomization often leads to reduced sample efficiency [11]. Alternatively, many works incorporate known physics principles or state structures into task design to approach task generalization. Examples include foot-placement dynamics [15], object invariance [16], granular media interactions [17], frequency domain parameterization [7], rigid body dynamics [18], and semi-structured Lagrangian dynamics models [19].
Some prior works have explored experience augmentation in RL, where original state-action trajectories are modified to enrich the training signal. For example, PlayVirtual [20] augments virtual trajectories using forward and backward dynamics in a latent space to enhance representation learning. Hindsight Experience Replay (HER)  [12, 21, 22, 23, 24] relabels failed trajectories with alternative goals to provide more informative learning signals. In the context of legged locomotion, symmetry-based augmentation has been applied to encourage symmetry-invariant behavior [25, 26]. While these methods leverage experience augmentation to improve policy learning, they do not explicitly tackle the challenge of task generalization.
Another relevant line of research is meta-RL [27, 28, 29, 30], which enables agents to quickly adapt to new tasks by leveraging shared structure across tasks encountered during meta-training. Memory-based meta-learners [29, 30], which encode task context through interaction history, have been leveraged for applications such as learning dynamics model priors for online adaptation [31] and training morphology-agnostic locomotion policies [32]. In LDM [33] and VariBAD Dream [34], experience augmentation is integrated into meta-RL by generating imagined tasks from the learned latent task space of a VariBAD [28] model to improve generalization. However, most meta-RL methods [27, 28, 29, 30, 34] focus on ID generalization, assuming test tasks are sampled from the same distribution as training tasks, and are not inherently equipped to handle OOD generalization. While LDM takes a step toward addressing this gap by generating imagined tasks through interpolation in the latent task space for OOD generalization, the resulting tasks may not accurately capture meaningful task structures, limiting generalization in complex environments. In this work, we contribute to addressing this challenge by integrating task-structured experience augmentation with memory-based meta-RL, enabling generalization to plausible OOD tasks.
3 Preliminaries
3.1 Partially observable Markov Decision Process and Meta-RL
@@
We model the problem as a Partially Observable Markov Decision Process (POMDP). We define a POMDP as the tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{O},\mathcal{A},p,r,\gamma,\rho_{0}\rangle,$ where $\mathcal{S}$ is the set of states, $\mathcal{O}$ is the set of observations, $\mathcal{A}$ is the set of actions, $p:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}_{\geq 0}$ is the state transition function, $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is the reward function, $\gamma$ is the discount factor, and $\rho_{0}:\mathcal{S}\to\mathbb{R}_{\geq 0}$ is the initial state distribution. In POMDPs, observations provide only partial information about the true underlying state, requiring the agent to infer the latent state. In our setup, this latent state corresponds to the unobservable task context, such as an unknown payload on the robot’s body, which remains fixed throughout each rollout. In conventional RL, the agent aims to learn a policy $\pi$ that maximizes the expected cumulative reward:
@@
where $\tau=(s_{0},a_{0},s_{1},\dots)$ represents a trajectory sampled from the POMDP $\mathcal{M}$ under policy $\pi$.
In this work, we frame the problem of learning across multiple tasks as a meta-RL problem, with the objective of optimizing over a distribution of tasks $p(\mathcal{T})$. In our setting, this corresponds to a distribution of POMDPs $\mathcal{M}(\mathcal{T})$. The goal is to learn a policy $\pi$ that maximizes the expected cumulative reward across tasks:
@@
3.2 Problem statement
@@
During training, only partial observations are available for $\mathcal{T}_{\text{ID}}$. In this setting, the agent must infer the latent task context implicitly based on past observations and actions.
4 Approach
4.1 Task-structured augmentation
To augment the training experience, our approach uses a set of transformation functions, denoted by $\mathcal{G}$, where each $g\in\mathcal{G}$ represents a transformation reflecting the underlying structure assumed to exist in the task space. Formally, let $\tau=\{(o_{t},a_{t},o_{t+1},r_{t})\}_{t=0}^{T}$ represent a trajectory collected by the current policy $\pi_{\theta_{k}}$ within the training set $\mathcal{T_{\text{ID}}}$. Each transformation $g=(g_{o},g_{a})\in\mathcal{G}$ consists of functions $g_{o}:\mathcal{O}\rightarrow\mathcal{O}$ and $g_{a}:\mathcal{A}\rightarrow\mathcal{A}$, which map an observation-action pair $(o,a)$ to a new pair $(o^{g},a^{g})=(g_{o}(o),g_{a}(a))$. Although transformations are applied at the observation level, we assume they are induced by consistent transformations at the underlying state level. For all $g\in\mathcal{G},\,s\in\mathcal{S},\,a\in\mathcal{A}$, it is assumed that the following properties hold:
@@ 4.1 Task-structured augmentation
    Transition probability invariance: The effect of actions on state transitions remains unchanged under transformation, i.e., $p(s_{t+1}^{g}\mid s_{t}^{g},a_{t}^{g})=p(s_{t+1}\mid s_{t},a_{t})$.
@@
    Initial state distribution invariance: Augmented episodes start from statistically equivalent conditions, i.e., $\rho_{0}(s^{g})=\rho_{0}(s)$.
@@
    Reward invariance: Transformed and original state-action pairs receive the same rewards in their respective tasks, i.e., $r(s_{t}^{g},a_{t}^{g})=r(s_{t},a_{t})$.
These properties ensure that the transformed trajectories are behaviorally equivalent to the original ones in their respective tasks, allowing us to reuse the advantage estimates and action probabilities from the original rollouts during policy updates (see Sec. 4.2). The resulting transformed trajectory $\tau_{g}=\{(o_{t}^{g},a_{t}^{g},o_{t+1}^{g},r_{t})\}_{t=0}^{T}$ represents the agent’s simulated experience within an augmented task $T\in\mathcal{T}_{\text{OOD}}^{\text{aug}}$, generated without real environment interactions.
An example of such transformations in $\mathcal{G}$ includes symmetry transformations, which have been used in RL [35] for manipulation tasks [22] and legged locomotion control [25, 26]. Specifically, for quadrupedal robots, symmetry transformations can simulate mirrored conditions (e.g., joint failures) across left-right and front-back body parts. Similarly, for humanoid robots, they can replicate mirrored conditions across left-right body parts. These transformations leverage the morphological symmetry inherent to legged robots and uphold the essential invariance properties (a)(c) [36].
4.2 Training with augmented data
We use Proximal Policy Optimization (PPO) [37] as the underlying RL algorithm. Following the method of Mittal et al. [25], we approach data augmentation in PPO by constructing augmented policies $\pi_{\theta_{k}}^{g}$ such that $\pi_{\theta_{k}}^{g}(a^{g}|s^{g})=\pi_{\theta_{k}}(a|s),\forall g\in\mathcal{G},s\in\mathcal{S},a\in\mathcal{A}.$ Based on properties (a)(c), it can be shown that the advantage function and state probability distribution are preserved under transformation, i.e., $A^{\pi_{\theta_{k}}^{g}}(s^{g},a^{g})=A^{\pi_{\theta_{k}}}(s,a)\text{, and }p_{\pi_{\theta_{k}}^{g}}(s^{g})=p_{\pi_{\theta_{k}}}(s)$ [25]. These invariances allow the policy update to incorporate augmented trajectories while retaining the action probabilities of the original data. For further details, we refer readers to [25].
@@ 4.2 Training with augmented data
This approach ensures balanced learning across both the original and augmented tasks by maintaining equal representation for each task during the policy update.
4.3 Memory augmentation
While experience augmentation can enhance generalization by increasing training diversity, the resulting variations, combined with the partially observable nature of the environment, may cause the agent to adopt over-conservative strategies as it struggles to infer the latent task context. To address this limitation, we extend augmentation to the agent’s memory by incorporating an RNN into the training framework, as illustrated in Fig. 1. The RNN, leveraging its context modeling capability, acts as an implicit task encoder. At each time step, the RNN updates its hidden state $h_{t}$ based on the observations, progressively capturing task-specific information accumulated so far. From this hidden state, the RNN generates a latent task embedding $z_{t}$, which conditions the policy, allowing it to adapt decisions based on the inferred context.